Which would you rather have: A job that pays $1000 for 30 days or a job that pays you a penny on the first day, 2 pennies on the second day, 4 pennies on the third, 8 pennies on the forth and continues each day for 30 days? Show mathematical evidence to support your answer.
Answer by htmentor(1343)

The amount of money you earn (in pennies) on the n-th day will be 2^(n-1).
The job that pays $1000 earns you 100000 pennies.
Is there a day (n<30) in which you start to earn more than 100000 pennies?
2^(n-1) = 100000 = 10^5
log10(2^(n-1)) = log10(10^5)
(n-1)log10(2) = 5
n = 5/log10(2) + 1
This gives n = 17.6. So on the 18th day, you will already start earning more than the $1000, not even counting all the money you've earned up to that point.
Take the penny-doubling job...